OpenStudy (samsan9):

which of the following could be an example of a function with the domain [a,infinity) and a range (-infinity,b) where a>0 and b>0?

OpenStudy (samsan9):

\[f(x)=\sqrt{x+a}-b\]

OpenStudy (samsan9):

\[\sqrt[3]{x+a}-b\]

OpenStudy (samsan9):

\[x=\sqrt[3]{(x-b)}+a\]

OpenStudy (samsan9):

or \[f(x)=-\sqrt{x-a}+b\]

OpenStudy (samsan9):

my bad it is supposed to be (-infinity,b]

OpenStudy (samsan9):

@goformit100

OpenStudy (samsan9):

@arithmetic555